# Quantum geometry from the Moyal product: quantum kinetic equation and non-linear response

### v1

@@ 1 Introduction
While the most famous kinetic equation is the semiclassical Boltzmann equation for phase space densities, capturing the quantum coherence of particles requires using and “diagonalizing” quantum kinetic equations which involve full density matrices. Although these equations can be formulated exactly in phase phase using the Moyal formalism [1], diagonalizing them exactly is usually not possible. However, this formalism is well suited to the case of small gradients, for which one can systematically expand the equations order by order in gradients, subsequently diagonalize them, and express observables using “single-band quantities.” Under this procedure, the quantum coherence effects, in particular the multiple-band effects, manifest in each individual band through “geometric terms”, which involve derivatives of the diagonalization matrix (and in turn eigenstates of all bands). At first order in the gradients, one recovers the usual Boltzmann’s equation with the proper velocity and force using this formalism [2, 3, 4, 5]. At that order, the real, momentum and mixed space components of the Berry curvature enter. Identifying some coefficients of phase space derivatives of the distribution function with phase space velocities, one can also recover the single particle equations of motion of the center of a wavepacket in the “wavepacket formalism” [6], which can be further formulated as symplectic mechanics [7, 8, 9, 10].
@@
Here, we push the Moyal product formalism to second order in the gradients for particle-conserving systems of non-interacting fermions without assuming translation symmetry. We show that the Berry curvature gains corrections at this order, and that the quantum metric tensor and a symmetric interband coherence tensor modify the kinetic equation and the observables. Our results are general and exact to second order in the gradients, and provide a unified framework that can be applied to many different problems including transport and dynamical response. Notably, using this systematic approach, we find in particular that the wavepacket formalism does not capture all contributions to response functions, and working up to second order also allows us to highlight, even at first order, an ambiguity in the definition of band diagonal quantities.
@@ 2.2 Moyal diagonalization
We can take the star-unitary operators that are used to diagonalize ${\sf H}$ to define $\tilde{F}\equiv U^{\dagger}\star{\sf F}\star U$. Although ${\sf H}$ is diagonalized by $U$, $\tilde{F}$ is generally not a diagonal matrix. However only the diagonal components of $\tilde{F}$ contribute to observables such as the total particle number and energy. For convenience, we separate the diagonal and off-diagonal components of $\tilde{F}$ labeling them $\tilde{f}$ and $\tilde{\mathcal{F}}$ respectively such that $\tilde{F}=\tilde{f}+\tilde{\mathcal{F}}$.
@@ 2.5 Kinetic equation
where $\Lambda_{t}\equiv-iU^{\dagger}\star\partial_{t}U$. For time-independent Hamiltonians and for a specific class of time-dependent Hamiltonians that can be expressed as ${\sf H}(t)=\sum_{n=1}^{N}P_{n}\star h_{n}(t)\star P_{n}$ where $\{P_{n}\}$ are time-independent projection operators and $\{h_{n}(t)\}$ are time-dependent scalar functions (c.f. appendix H), $\Lambda_{t}=0$. In this case, if $\tilde{F}$ is diagonal at some past time, it will remain diagonal at later times. However, here, we assume $\Lambda_{t}\neq 0$ and proceed assuming $\tilde{F}$ is generally non-diagonal. eq. 30 can be separated into two coupled kinetic equations for the diagonal ($\tilde{f}$) and off-diagonal ($\tilde{\mathcal{F}}$) components of $\tilde{F}$:
$i\hbar\partial_{t}\tilde{f}=$ $[\tilde{h}+\hbar A_{t}\,\overset{\star}{,}\,\tilde{f}]+\hbar\operatorname{diag}([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]),$ (31)
$i\hbar\partial_{t}\tilde{\mathcal{F}}=$ $[\tilde{h}+\hbar\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]-\hbar\operatorname{diag}([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}])$ (32)
@@ 2.6 Equilibrium distribution function $f_{\textrm{eq}}$
where $n_{F}=n_{F}(h)$ We provide the details of the calculation leading to (38) in appendix E. $f_{\rm eq}$ in (38) solves (33) with $\partial_{t}f_{\rm eq}=0$ (equilibrium) and $\Lambda_{t}=0$.
@@ 2.8 Electric current density
The electric current density can be obtained by introducing an external $U(1)$ gauge field $\mathcal{A}_{i}(x)$ and taking the derivative of the Hamiltonian with respect to the gauge field, i.e. $J_{i}(x)\equiv-\delta\langle{\rm\hat{H}}\rangle/\delta\mathcal{A}_{i}(x)$ (see appendix F). What we find to second order in $\hbar$ is
@@
Assuming the off-diagonal components $\tilde{\mathcal{F}}$ are negligible, we can express the current density using the diagonalized Hamiltonian $h$ and the diagonal components of the distribution function $f$. The details of this calculation are provided in appendix I. We find that the current density can be expressed as
@@
The contributions to current density from the terms $b^{a}_{ij},b^{m}_{ij;k}$ are divergenceless, so they correspond to bound currents. $c_{ij}$, which appears in $b_{ij;k}^{m}$, has been referred to as the “band Drude weight” in existing literature [24] and its momentum integral over occupied states has been shown to be the correction to the free-electron Drude weight coming from the periodic potential of a crystal [25, 26].
@@ 2.9 Uniform magnetic field
A uniform magnetic field can be introduced in our formalism relatively simply by changing variables from the phase-space coordinates $(x,p)$ to $(x,\pi)$ where $\pi_{i}=p_{i}+e\mathcal{A}_{i}$ is kinematic momentum, where $\mathcal{A}_{i}$ is the external vector potential, as before. Under this change of variables, the Moyal product becomes [27, 4, 28]
$\star_{B}=$ $\exp\left(\frac{i\hbar}{2}({\mathchoice{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\displaystyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\displaystyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\textstyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\textstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptstyle\vec{}\mkern 4.0mu$}\cr\kern-3.01389pt\cr$\scriptstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptscriptstyle\vec{}\mkern 4.0mu$}\cr\kern-2.15277pt\cr$\scriptscriptstyle\nabla$\cr}}}}_{x}\cdot\vec{\nabla}_{\pi}-{\mathchoice{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\displaystyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\displaystyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\textstyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\textstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptstyle\vec{}\mkern 4.0mu$}\cr\kern-3.01389pt\cr$\scriptstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptscriptstyle\vec{}\mkern 4.0mu$}\cr\kern-2.15277pt\cr$\scriptscriptstyle\nabla$\cr}}}}_{\pi}\cdot\vec{\nabla}_{x})-\frac{ie\hbar}{2}\mathfrak{F}_{ij}{\mathchoice{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\displaystyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\displaystyle\partial$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\textstyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\textstyle\partial$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptstyle\vec{}\mkern 4.0mu$}\cr\kern-3.01389pt\cr$\scriptstyle\partial$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptscriptstyle\vec{}\mkern 4.0mu$}\cr\kern-2.15277pt\cr$\scriptscriptstyle\partial$\cr}}}}_{\pi_{i}}\vec{\partial}_{\pi_{j}}\right)$ (79)
@@
In appendix J, we apply this magnetic Moyal product to the equilibrium current, and find that the divergence of the latter is non-vanishing at a Weyl node due to the chiral anomaly effect [7, 29]. A more detailed study of the magnetic field contributions to second order in $\hbar$ is left for the future.
@@ 3 Application: Transport driven by electric fields
where ${\sf H}_{0}(p)$ is a $N\times N$ Hamiltonian that is generally non diagonal and $-e<0$ is the electric charge of the fermion. As mentioned before, at zeroth order in $\hbar$, the diagonalization equation reduces to a matrix diagonalization problem, so the columns of $U^{(0)}$ are the eigenvectors of ${\sf H}_{0}(p)$. If the eigenvector and eigenvalues of ${\sf H}_{0}(p)$ are given by $\left\{|u_{n}(p)\rangle\right\}$ and $\left\{\varepsilon_{n}(p)\right\}$,
@@
Here, ${\sf P}_{n}$ is the projection operator to zeroth order in $\hbar$ given by ${\sf P}_{n}\equiv P_{n}^{(0)}=|u_{n}\rangle\langle u_{n}|$. As mentioned in the introduction, $t_{n,ij}$ in existing literature is referred to as either the “Berry connection polarizability” [15, 19] or as the “band-normalized quantum metric” [20, 18]. We want to emphasize that in general this quantity is not a band-geometric quantity since it is a multi-band quantity as its calculation relies on the knowledge of all bands. Only in the case of a two-band model can it truly be considered a band-geometric quantity since the energy denominator in the summation is simply replaced by the energy difference of the two bands.
@@ 3.1 Equilibrium current
The first two terms in the first line are expressed using the Berry connection which is gauge dependent. However, as mentioned before, the current itself is gauge-invariant. Indeed the derivative $\partial_{x_{j}}$ acts on $f_{n}$ to produce $-f_{n}^{\prime}e\partial_{x_{i}}V$. Then, by integration by parts, these two terms together can be rewritten as $-(e^{2}/\hbar)\sum_{n=1}^{N}\int_{k}f_{n}{\sf\Omega}_{n,ij}\partial_{x_{j}}V$ which, of course, is the Hall current. As mentioned earlier, this expression is valid when $\{\varepsilon_{n}\}$ are non-degenerate. For the case of a Weyl fermion, discussed in appendix J, the Berry connection ${\sf A}$ is ill-defined at the Weyl point, so the magnetization $\mathcal{M}_{ij}$ cannot be defined.
@@ 3.2.1 Kinetic equation
where we assumed $\tilde{\mathcal{F}}$ vanishes in the absence of an external potential. If we define the Fourier transform ~$\tilde{\mathcal{F}}(x,k,t)=\int_{q,\omega}\widehat{\tilde{\mathcal{F}}}_{nm}(q,k,\omega)e^{iq\cdot x-i\omega t}$, we find to first order in small $q$ (the expansion in $q$ is equivalent to the expansion in spatial gradients)
(~^$(\widehat{\tilde{\mathcal{F}}})_{nm}=$ $-ie\hbar\frac{({\sf\Lambda}_{i})_{nm}}{\varepsilon_{n}-\varepsilon_{m}}\frac{n_{F,n}-n_{F,m}}{\hbar\omega-\varepsilon_{n}+\varepsilon_{m}}\omega q_{i}\widehat{V}(q,\omega)$
@@
where $n_{F,n}=n_{F}(\varepsilon_{n})$. At this order, if $\hbar\omega,eV\ll\varepsilon_{n}-\varepsilon_{m}$, we see that (~^$(\widehat{\tilde{\mathcal{F}}})_{nm}\sim\hbar\omega e\widehat{V}/(\varepsilon_{n}-\varepsilon_{m})^{2}\ll 1$. Therefore, in the low-frequency limit $\omega\ll\omega_{\textrm{gap}}$ where $\hbar\omega_{\textrm{gap}}$ is the smallest of the band gaps neighboring the Fermi energy, the contribution from $\tilde{\mathcal{F}}$ is negligible. Physically, this simply means the off-diagonal components can be ignored if we consider frequencies that are sufficiently small such that the external potential cannot induce interband transitions. This is the justification for the assumption that we made earlier that ~$\tilde{\mathcal{F}}=0$.
@@ 3.2.2 Definition of the relaxation time approximation
The details of the derivation are described in appendix K.
@@ 3.2.3 Solutions to the kinetic equation and current (expansion in small potential)
$-e^{2}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{n_{F}^{\prime}v_{i}{\sf t}_{jl}q_{i}q^{\prime\prime}_{j}q^{\prime\prime}_{l}}{\hbar(\omega-\vec{v}\cdot\vec{q}+i\tau^{-1})}\widehat{V}(q^{\prime},\omega^{\prime})\widehat{V}(q^{\prime\prime},\omega^{\prime\prime})-\frac{\omega-\vec{v}\cdot\vec{q}}{\omega-\vec{v}\cdot\vec{q}+i\tau^{-1}}\widehat{f}^{r}_{e=2}(q,k,\omega)+\mathcal{O}(q^{3})$
@@
where we defined $\langle d\rangle_{\textrm{FS}}\equiv\sum_{n=1}^{N}\int_{k}n_{F,n}^{\prime}d_{n}$ as an integral over the Fermi surface. The first two terms are the familiar Drude term ($\frac{e^{2}}{\hbar}\frac{\hbar\tau\langle v_{i}v_{j}\rangle_{\textrm{FS}}}{1-i\omega\tau}$) and Hall effect term ($-\frac{e^{2}}{\hbar}\sum_{n=1}^{N}\int_{k}n_{F,n}{\sf\Omega}_{n,ij}$) respectively. The other terms describe the electric current linear response to a non-uniform electric field. Notice that all the terms in the last line are divergenceless terms that describe edge currents and they are also multi-band quantities (meaning their calculation requires knowledge of the wavefunctions of multiple bands) in contrast to the terms in the first line which are all single-band quantities (whose calculation only requires knowledge of the wavefunction of a single band). Also note that a few terms in (122) remain nonzero even when the Fermi energy is in a band gap at $T=0$. This does not however contradict what we expect for a band insulator since those terms which are nonzero are divergenceless currents and conserve the local charge distribution.
@@
Note that the second term of eq. 123, i.e. $-\frac{e^{2}}{\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{\tau\langle v_{l}{\sf\Omega}_{ij}\rangle_{\textrm{FS}}}{1-i\omega^{\prime\prime}\tau}$, which is a transverse contribution to the current, is exactly the nonlinear Hall effect first proposed by Sodemann and Fu [21] that originates from the Berry curvature dipole.
If the Fermi energy is in a band gap at $T=0$, all terms except the last vanish, but unlike in the case of linear electrical response this finite term is not divergenceless. This occurs because the RTA violates local charge conservation in the system. The RTA couples the system to an external bath using a constant scattering rate, so even if the Fermi energy is in the band gap, a “divergenceful” current can still be generated. A pratical remedy for this issue is to take the collisionless limit $\tau^{-1}\rightarrow 0$ when the Fermi energy is in the band gap.
A general expression for second-order conductivity that ignored interband transitions was also studied in Ref. [18] by Kaplan et al. in which they used the RTA. Because their choice of $f^{r}$ in the RTA differs from ours and is $f^{r}=n_{F}(\varepsilon)$, which differs from ours, only the first line of eq. 123 here should be compared with the results of Ref. [18]. [Indeed the last line of eq. 123 comes from $f^{r}_{e=2}$ of eq. 117, which does not exist in Ref. [18].] Moreover, Kaplan et al. express their results as an integral over the Fermi sea, so we must we write our $\langle\cdots\rangle_{\textrm{FS}}$ terms by integrating by parts to compare them with their results. What we find is that only the first and last terms of the first line in eq. 123 match the results of Ref. [18] exactly. Contributions proportional to the second and third term are also present in the results of Ref. [18], but we find the numerical prefactors are different.

### v2

@@ 1 Introduction
While the most famous kinetic equation is the semiclassical Boltzmann equation for phase space densities, capturing the quantum coherence of particles requires using and “diagonalizing” quantum kinetic equations which involve full density matrices. Although these equations can be formulated exactly in phase phase using the Moyal formalism [20], diagonalizing them exactly is usually not possible. However, this formalism is well suited to the case of small gradients, for which one can systematically expand the equations order by order in gradients, subsequently diagonalize them, and express observables using “single-band quantities.” Under this procedure, the quantum coherence effects, in particular the multiple-band effects, manifest in each individual band through “geometric terms”, which involve derivatives of the diagonalization matrix (and in turn eigenstates of all bands). At first order in the gradients, one recovers the usual Boltzmann’s equation with the proper velocity and force using this formalism [6, 10, 11, 40, 29, 18]. At that order, the real, momentum and mixed space components of the Berry curvature enter. Identifying some coefficients of phase space derivatives of the distribution function with phase space velocities, one can also recover the single particle equations of motion of the center of a wavepacket in the “wavepacket formalism” [41], which can be further formulated as symplectic mechanics [32, 33, 5, 12].
@@
Here, we push the Moyal product formalism to second order in the gradients for particle-conserving systems of non-interacting fermions without assuming translation symmetry. Our approach and results are closely related to the pioneering works of Blount[2, 3, 4], but extend and systematize the procedure to greater generality and to higher accuracy. We show that the Berry curvature gains corrections at this order, and that the quantum metric tensor and a symmetric interband coherence tensor modify the kinetic equation and the observables. Our results are general and exact to second order in the gradients, and provide a unified framework that can be applied to many different problems including transport and dynamical response. Notably, using this systematic approach, we find in particular that the wavepacket formalism does not capture all contributions to response functions, and working up to second order also allows us to highlight, even at first order, an ambiguity in the definition of band diagonal quantities.
@@ 2.2 Moyal diagonalization
We can take the star-unitary operators that are used to diagonalize $\sf H$} to define $\tilde{F}\equiv U^{\dagger}\star{\sf F}\star U$. Although $\sf H$} is diagonalized by $U$, $\tilde{F}$ is generally not a diagonal matrix. However only the diagonal components of $\tilde{F}$ contribute to observables such as the total particle number and energy. For convenience, we separate the diagonal and off-diagonal components of $\tilde{F}$ labeling them $\tilde{f}$ and $\tilde{\mathcal{F}}$ respectively such that $\tilde{F}=\tilde{f}$$\tilde{\mathcal{F}}$.
@@ 2.5 Kinetic equation
where $\Lambda_{t}\equiv-iU^{\dagger}\star\partial_{t}U$. For time-independent Hamiltonians and for a specific class of time-dependent Hamiltonians that can be expressed as ${\sf H}(t)=\sum_{n=1}^{N}P_{n}\star h_{n}(t)\star P_{n}$ where $\{P_{n}\}$ are time-independent projection operators and $\{h_{n}(t)\}$ are time-dependent scalar functions (c.f. appendix I), $\Lambda_{t}=0$. In this case, if $\tilde{F}$ is diagonal at some past time, it will remain diagonal at later times. However, here, we assume $\Lambda_{t}\neq 0$ and proceed assuming $\tilde{F}$ is generally non-diagonal. eq. 30 can be separated into two coupled kinetic equations for the diagonal ($\tilde{f}$) and off-diagonal ($\tilde{\mathcal{F}}$) components of $\tilde{F}$:
[…]
@@ 2.6 Equilibrium distribution function $f_{\textrm{eq}}$
where $n_{F}=n_{F}(h)$. We provide the details of the calculation leading to (38) in appendix E. $f_{\rm eq}$ in (38) solves (33) with $\partial_{t}f_{\rm eq}=0$ (equilibrium) and $\Lambda_{t}=0$.
The equilibrium distribution of an insulator can be written as the divergence of the polarization. In App. F, we extract the polarization from (38) and find partial agreement with the results of [42], which uses the semiclassical method.
@@ 2.8 Electric current density
The electric current density can be obtained by introducing an external $U(1)$ gauge field $\mathcal{A}_{i}(x)$ and taking the derivative of the Hamiltonian with respect to the gauge field, i.e. $J_{i}(x)\equiv-\delta\langle{\rm\hat{H}}\rangle/\delta\mathcal{A}_{i}(x)$ (see appendix G). What we find to second order in $\hbar$ is
@@
Assuming the off-diagonal components $\tilde{\mathcal{F}}$ are negligible, we can express the current density using the diagonalized Hamiltonian $h$ and the diagonal components of the distribution function $f$. The details of this calculation are provided in appendix J. We find that the current density can be expressed as
@@
The contributions to current density from the terms $b^{a}_{ij},b^{m}_{ij;k}$ are divergenceless, so they correspond to bound currents. $c_{ij}$, which appears in $b_{ij;k}^{m}$, has been referred to as the “band Drude weight” in existing literature [14] and its momentum integral over occupied [25, 26].
@@ 2.9 Uniform magnetic field
A uniform magnetic field can be introduced in our formalism relatively simply by changing variables from the phase-space coordinates $(x,p)$ to $(x,\pi)$ where $\pi_{i}=p_{i}+e\mathcal{A}_{i}$ is kinematic momentum, where $\mathcal{A}_{i}$ is the external vector potential, as before. Under this change of variables, the Moyal product becomes [22, 40, 45, 44]
[…]
@@
In appendix K, we apply this magnetic Moyal product to the equilibrium current, and find that the divergence of the latter is non-vanishing at a Weyl node due to the chiral anomaly effect [32, 34]. In appendix N, we derive the band-diagonal gauge-invariant Hamiltonian, showing that in a uniform magnetic field, it only depends on the kinetic momentum. A more detailed study of the transport under the uniform magnetic field to the second order in $\hbar$ is left for the future.
@@ 3 Application: Transport driven by electric fields
where ${\sf H}_{0}(p)$ is a $N\times N$ Hamiltonian that is generally non diagonal and e<0-e<0 is the electric charge of the fermion. As mentioned before, at zeroth order in $\hbar$, the diagonalization equation reduces to a matrix diagonalization problem, so the columns of $U^{(0)}$ are the eigenvectors of ${\sf H}_{0}(p)$. If the eigenvector and eigenvalues of ${\sf H}_{0}(p)$ are given by $\left\{|u_{n}(p)\rangle\right\}$ and $\left\{\varepsilon_{n}(p)\right\}$,
@@
Here, ${\sf P}_{n}$ is the projection operator to zeroth order in $\hbar$ given by ${\sf P}_{n}\equiv P_{n}^{(0)}=|u_{n}\rangle\langle u_{n}|$. We find (95) is in agreement with the “positional shift” in [9]. As mentioned in the introduction, $t_{n,ij}$ in existing literature is referred to as either the “Berry connection polarizability” [9, 16] or as the “band-normalized quantum metric” [39, 15]. We want to emphasize that in general this quantity is not a band-geometric quantity since it is a multi-band quantity as its calculation relies on the knowledge of all bands. Only in the case of a two-band model can it truly be considered a band-geometric quantity since the energy denominator in the summation is simply replaced by the energy difference of the two bands.
@@ 3.1 Equilibrium current
The first two terms in the first line are expressed using the Berry connection which is gauge dependent. However, as mentioned before, the current itself is gauge-invariant. Indeed the derivative $\partial_{x_{j}}$ acts on $f_{n}$ to produce $f_{n}^{\prime}e\partial_{x_{i}}V$. Then, by integration by parts, these two terms together can be rewritten as $(e^{2}/\hbar)\sum_{n=1}^{N}\int_{k}f_{n}{\sf\Omega}_{n,ij}\partial_{x_{j}}V$ which, of course, is the anomalous Hall current. Moreover, by including the orbital magnetization (the third term in the first line), the first-order terms of (111) agree with the semiclassical results in [41, 36].
As mentioned earlier, this expression is valid when $\{\varepsilon_{n}\}$ are non-degenerate. For the case of a Weyl fermion, discussed in appendix K, the Berry connection $\sf A$} is ill-defined at the Weyl point, so the magnetization $\mathcal{M}_{ij}$ cannot be defined.
@@ 3.2.1 Kinetic equation
where we assumed $\tilde{\mathcal{F}}$ vanishes in the absence of an external potential. If we define the Fourier transform $\tilde{\mathcal{F}}(x,k,t)=\int_{q,\omega}\widehat{\tilde{\mathcal{F}}}_{nm}(q,k,\omega)e^{iq\cdot x-i\omega t}$, we find to first order in small $q$ (the expansion in $q$ is equivalent to the expansion in spatial gradients)
$(\widehat{\tilde{\mathcal{F}}})_{nm}=$ $-ie\hbar\frac{({\sf\Lambda}_{i})_{nm}}{\varepsilon_{n}-\varepsilon_{m}}\frac{n_{F,n}-n_{F,m}}{\hbar\omega-\varepsilon_{n}+\varepsilon_{m}}\omega q_{i}\widehat{V}(q,\omega)$
@@
where $n_{F,n}=n_{F}(\varepsilon_{n})$. At this order, if $\hbar\omega,eV\ll\varepsilon_{n}-\varepsilon_{m}$, we see that $(\widehat{\tilde{\mathcal{F}}})_{nm}\sim\hbar\omega e\widehat{V}/(\varepsilon_{n}-\varepsilon_{m})^{2}\ll 1$. Therefore, in the low-frequency limit $\omega\ll\omega_{\textrm{gap}}$ where $\hbar\omega_{\textrm{gap}}$ is the smallest of the band gaps neighboring the Fermi energy, the contribution from $\tilde{\mathcal{F}}$ is negligible. Physically, this simply means the off-diagonal components can be ignored if we consider frequencies that are sufficiently small such that the external potential cannot induce interband transitions. This is the justification for the assumption that we made earlier that $\tilde{\mathcal{F}}=0$.
@@ 3.2.2 Definition of the relaxation time approximation
The details of the derivation are described in appendix L.
@@ 3.2.3 Solutions to the kinetic equation and current (expansion in small potential)
$+e^{2}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{n_{F}^{\prime}v_{i}{\sf t}_{jl}q^{\prime}_{i}q^{\prime\prime}_{j}q^{\prime\prime}_{l}}{\hbar(\omega-\vec{v}\cdot\vec{q}+i\tau^{-1})}\widehat{V}(q^{\prime},\omega^{\prime})\widehat{V}(q^{\prime\prime},\omega^{\prime\prime})-\frac{\omega-\vec{v}\cdot\vec{q}}{\omega-\vec{v}\cdot\vec{q}+i\tau^{-1}}\widehat{f}^{r}_{e=2}(q,k,\omega)+\mathcal{O}(q^{3})$
@@
where we defined $\langle d\rangle_{\textrm{FS}}\equiv\sum_{n=1}^{N}\int_{k}n_{F,n}^{\prime}d_{n}$ as an integral over the Fermi surface. The first two terms are the familiar Drude term ($\frac{e^{2}}{\hbar}\frac{\hbar\tau\langle v_{i}v_{j}\rangle_{\textrm{FS}}}{1-i\omega\tau}$) and Hall effect term ($\frac{e^{2}}{\hbar}\sum_{n=1}^{N}\int_{k}n_{F,n}{\sf\Omega}_{n,ij}$) respectively. The other terms describe the electric current linear response to a non-uniform electric field. Notice that all the terms in the last line are divergenceless terms that describe edge currents and they are also multi-band quantities (meaning their calculation requires knowledge of the wavefunctions of multiple bands) in contrast to the terms in the first line which are all single-band quantities (whose calculation only requires knowledge of the wavefunction of a single band). Also note that a few terms in (122) remain nonzero even when the Fermi energy is in a band gap at $T=0$. This does not however contradict what we expect for a band insulator since those terms which are nonzero are divergenceless currents and conserve the local charge distribution.
@@
Note that the second term of eq. 123, i.e. $\frac{e^{2}}{\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{\tau\langle v_{l}{\sf\Omega}_{ij}\rangle_{\textrm{FS}}}{1-i\omega^{\prime\prime}\tau}$, which is a transverse contribution to the current, is exactly the nonlinear Hall effect first proposed by Sodemann and Fu [31] that originates from the Berry curvature dipole.
If the Fermi energy is in a band gap at $T=0$, all terms except the last vanish, but unlike in the case of linear electrical response this finite term is not divergenceless. This occurs because the RTA violates local charge conservation in the system. The RTA couples the system to an external bath using a constant scattering rate, so even if the Fermi energy is in the band gap, a “divergenceful” current can still be generated. A practical remedy for this issue is to take the collisionless limit $\tau^{-1}\rightarrow 0$ when the Fermi energy is in the band gap.
A general expression for second-order conductivity that ignored interband transitions was also studied in Ref. [15] by Kaplan et al. in which they chose $f^{r}=n_{F}(\varepsilon)$. If we use this choice of $f^{r}$ in our formalism for $\widehat{J}_{i,e=3}$, our result is modified to include only the first line of eq. 123, which should be directly comparable to the results of Ref. [15]. What we find is that only the first and last terms of the first line in eq. 123 match the results of Ref. [15] exactly. Contributions proportional to the second and third term are also present in the results of Ref. [15], but we find the numerical prefactors are different. Our results also extend those of Ref. [9] [^4] The nonlinear Hall effect in [9] can be seen as the transverse part of the nonlinear current that is proportional to ${\sf t}_{ij}$, $\widehat{J}_{i,e=3}\sim\frac{1}{2}e^{3}\int\langle v_{i}{\sf t}_{jl}-v_{j}{\sf t}_{il}\rangle_{\mathrm{FS}}q^{\prime}_{j}q^{\prime\prime}_{l}\widehat{V}(q^{\prime})\widehat{V}(q^{\prime\prime})$.